# Quantum Complexity of Nonlocal Field Theories

### v1

@@ 2.1 Volume complexity (CV)
where $u_{b}$ is the boundary UV-cutoff, the simplified extremal volume leads to complexity given by,
@@
where, $c$ is the central charge $c=\frac{N^{2}}{4}$ in large $N$ limit and $u_{b}=\frac{1}{\epsilon}$.
@@ 2.2 Subregion volume complexity
In this section we will study the subregion complexity of the dipole deformed theory, which lives on subregion along the noncompact boundary directions $\frac{-W}{2}<y,z<\frac{W}{2}$ and $\frac{-l}{2}<x<\frac{l}{2}$ with $W\rightarrow\infty$. The subregion complexity will be given by the maximal volume of co-dimension one spacelike surface bounded by the co-dimension two Ryu-Takanayagi(RT) surface, homologous to the subregion Alishahiha:2015rta . The metric on co-dimension two surface,
@@ 2.3 Action Complexity
where, $\lambda$ is t’Hooft coupling. The leading term, displays a volume law scaling and is independent of the dipole-deformation parameter $a$. Next, notice that the full action complexity, coincides with the pure AdS results which matches the volume complexity upto a numerical factor, when the the t’Hooft coupling is one. This is a unique feature of DDSYM theory that the complexity has an explicit dependence on the ’t-Hooft coupling parameter. Perhaps this points to the validity of holographic theory even for finite t’Hooft coupling, i.e. beyond the naive supergravity regime, in contrast to the local SYM-AdS duality case. Asymptotically (large $u_{b}$),
@@
    The strong nonlocality regime result (64) is worth considering for a moment. For complexity to be positive definite, the t’Hooft coupling $\lambda$ should satisfy the (upper) bound $\lambda^{2}<1+\ln{3}$. We speculate that this upper bound signals the breakdown of the UV-completeness of the DDSYM theory analogous to the LST case Katoch:2022hdf .
@@ 3 Noncommutative SYM theory: Anisotropic case
The impact of noncommutativity on the dual field theory is profound. Two-point correlation functions exhibit a characteristic UV/IR mixing — high energy (UV) behavior becomes entangled with long-distance (IR) physics, a signature of the underlying non-local interactions. Wilson loops, which in SYM serve as order parameters for confinement and phase transitions, acquire a dipole-like structure reflecting the noncommutative smearing of charge distributions. Furthermore, recent extensions of these analyses to quantum information theoretic measures, such as holographic entanglement entropy and computational complexity Karczmarek:2013xxa , reveal modified scaling laws and phase structures.
@@ 3.1 Volume complexity for the anisotropic case
$V(t_{*})$ $=\int d^{9}\sigma\penalty 10000\ e^{-2(\phi-\phi_{0})}\sqrt{G}=R^{9}L^{3}\omega_{5}\int_{u_{*}}^{u_{b}}du\,u^{2}\sqrt{1-u^{4}t^{\prime}(u)^{2}},$ (68)
It can easily be inferred from the inspection of (68) that the global maxima occurs when $t^{\prime}(u)=0$. implying that the volume complexity is
$\mathcal{C}_{V}(t_{*})$ $=\frac{L^{3}\omega_{5}}{3G_{N}^{(10)}}u_{b}^{3}$ (69)
Again, similar to the dipole case, the volume complexity is independent of noncommutative parameter $a$, because the deformation factors from the metric coefficients of the two noncommutative directions, and th edeformation factors arising from the dilaton profile cancels each other, a dimensional accident or conspiracy. In terms of the field theory picture, the loss of degrees of freedom due to noncommutativty in the UV is exactly compensated by the usual addition of degrees of freedom as one RG-flows from IR to the UV. As we will shortly see this dimensional conspiracy is evaded when we consider the fully isotropic NCSYM theory.
@@ 3.2 Subregion Volume complexity (CV) for the anisotropic case
To compute subregion complexity we focus our attention on the portion of the maximal volume slice which is contained within the Ryu-Takayanagi (RT) surface and is also homologous to the boundary subregion of interest given by the following range $\frac{-l}{2}<x<\frac{l}{2},\ \frac{-W}{2}<y,z<\frac{W}{2}$ with $W\to\infty$. In this case the embedding of the codim-2 surface is specified by the following geometry
@@
Since the above integration is not exactly solvable, we take two different regimes, namely Near AdS and far from AdS to obtain $l$ in terms of $u_{*}$ which will be useful to obtain Complexity in terms of sub region length in later subsections.
@@
Similar to near AdS case, we can also find $l$ in terms of $u_{*}$ in the Far from AdS regime as,
@@
The above integral is not solvable, so we plot the numerical integration results as shown in figure 2 for small values of the noncommutative deformation parameter (for large values of the deformation the numerics are not quite reliable or stable). The value of the critical subregion size (length) for three different values of $a$ extracted from the plots (in units of AdS radius) are as displayed in the table below. We also provide analytical estimates of the critical subregion size for comparison.
@@ 3.3 Action complexity (CA) for the anisotropic case (10 D)
The low energy limit of Type IIB string theory reduces to the following SUGRA bulk action with Neveu-Schwartz, Ramond-Ramond and Chern-Simmons sector outlined below
@@
  • Einstein-Hilbert term  [^7] Where we have used $\frac{\omega_{5}}{2\kappa^{2}_{10}}=\frac{\omega_{5}}{16\pi G_{N}^{(10)}}=\frac{1}{16\pi G_{N}^{(5)}}$
@@
The arguments of logarithm have been rendered dimensionless by rewriting them in the units of AdS radius. Moreover, $l_{ct}$ has been chosen to be one in the units of AdS radius without tempering the known universal result for the local field theory dual to AdS bulk. It can be seen that for the anisotropic case of noncommutative field theory, the action complexity is positive only in the range $\ln a<\frac{8}{3}$. Conversely positivity of action complexity imposes an upper bound on the noncommutativity parameter $a<\exp{\left(8/3\right)}$. This feature of action complexity is parallels with the action complexity of the DDSYM theory. However, unlike the DDSYM theory, but instead just like the local SYM field theory, the UV divergence scales with volume $L^{3}$! We postpone our speculations about this counterintuitive result till Sec. 5.
@@ 4 Noncommutative SYM theory: Isotropic case
In the previous two cases of DDSYM and anisotropic NCSYM, the deformation had a preferred direction and the full volume complexity turned out to be independent of the deformation parameter, i.e. turned out to be identical to undeformed SYM (pure AdS). We suspect this is a dimensional accident, i.e. artifact of the fact that there are only two spatial directions along which the deformation is turned on, and if the the deformation democratically involves all spatial and temporal directions, the full volume complexity will exhibit features distinct from the undeformed theory. To confirm our conjecture, in this section, we consider the Maldacena-Russo solution of a D3 brane in a constant NS $B$ field background given as Lin:2000ps ,
@@ 4.2 Action complexity (CA) for the isotropic case (5D)
Here we, as we did in the previous sections, in the end we have rescaled every argument of the logarithm in the units of AdS radius. Moreover, we have also set $l_{ct}$ to be one in the AdS radius scale without tempering the universal piece coming from the known pure AdS result. It is evident that the overall action complexity for the noncommutative case is negative unless the noncommutativity parameter satisfies the bound $\ln\left(\frac{a^{64}r_{b}^{58}}{5^{6}}\right)<4$ or $\left(\frac{a^{64}r_{b}^{58}}{5^{6}}\right)<e^{4}$. Moreover, the leading order scaling with the noncommutative coupling is the product of quadratic and logarithmic scaling with the nonlocal coupling parameter. It is however clear that due to the involvement of the UV scale the complexity cannot be made positive in any range of values in the coupling space. This is hinting at the theory being UV incomplete, this might imply that the theory has no ground state, or it has UV instability. Also, the action complexity in the noncommutative case scales with the fifth power of the UV cutoff, unlike the local field theory that is expected to scale cubically in the UV cutoff scale. This power law departure from the extensive scaling is interpreted as the effect of nonlocality on the action complexity.
@@ 5 Discussions and Outlook
Finally, we discuss few avenues for future work. The most natural and immediate extension would be to include the effects of finite temperature and finite charge/chemical potential in complexity for nonlocal field theories using the holographic charged black hole/brane SUGRA backgrounds instead. This is always an interesting exercise to carry out because the dual theory is nonlocal and it has the potential to shed light on novel physical effects associated purely with finite temperatures which might not appear in the local theories or for which we might not have any intuition from local theories. A second avenue to pursue is the subregion complexity for the isotropic NCSYM case for which the numerical estimates obtained using Mathematica or Python were unstable and unreliable for even small or moderate values of the deformation parameter $a$ - the numerical errors rapidly escalate to large values in course of the numerical solution of the RT curve. The numerics do not reproduce the pure AdS quadratic scaling of holographic entanglement entropy itself, let alone subregion complexity. It would be desirable to come up with a numerical scheme where the errors could be brought under control for moderate to large values of the deformation parameter $a$ and tie up this loose end in our present work. Another instructive, although perhaps less interesting avenue to explore will be to compute the subregion action complexity for all the cases considered in this paper. Finally, casting a wider net, one can investigate the holographic complexity (in addition to entanglement entropy) properties of more general nonlocal field theories dual to string backgrounds which are not asymptotically AdS, e.g. the nonlocal (C)FTs dual to null boundary holography Ferko:2025elh ; Dei:2025ryd ; Chakraborty:2025rfm . These are prototypes of celestial holography, and will help us explore holography/quantum gravity in backgrounds beyond the asymptotically AdS paradigm.
@@ Acknowledgements
The work of SP and SR is supported by the CRG grant of the Anusandhan National Research Foundation (ANRF), Department of Science and Technology (DST), CRG/2023/001120 (“Many facets of complexity: From chaos to thermalization”). The work of SR is also supported by the IIT Hyderabad seed grant SG/IITH/F171/2016-17/SG-47 as well the IITH funds RDF/IITH/F171/SR. The work of GK is supported by IIT Indore Sponsored Research Project no. CRG/2023/000904 (“From Modular Hamiltonian to Operator Algebra in Gauge/Gravity Dictionary”).
@@ Appendix A Entanglement entropy results for dipole deformed SYM
As the integral involved in area(17) can not be evaluated analytically we proceed numerical and study the behavior of area with subregion length (18) in figure 3. This shows a phase transition in entanglement entropy. This transition has been already observed in Karczmarek:2013xxa for a near boundary (large $u$) RT surface. Here we have done a full numerical analysis and observed that as subregion length increase the entanglement entropy (area) becomes independent of subregion length as well as the deformation parameter. In the near AdS region when $au<<1$, and the leading term in entanglement entropy is same as that of pure AdS. For small sub-region length degree of freedom inside the strip will be entangled to the degree of freedom outside the strip and the entanglement entropy will be proportional to the subregion length $l$. A limiting analysis of the area agrees with this interpretation.
@@
For near AdS region $(au<<1)$:
@@
Far from AdS region $(au>>1)$: The entanglement entropy is given by,

### v2

@@ 2.1 Volume complexity (CV)
where $u_{b}$ is the boundary UV-cutoff and the $W$ is the boundary IR cut off (the limits of integration for $W/2\leq x,y,z\leq W/2$). The simplified maximal volume determines the complexity,
@@
where, $c$ is the central charge $c=\frac{N^{2}}{4}$ in large $N$ limit and we have introduced the ”lattice spacing” regulator in the boundary $\epsilon$ via the bulk IR cutoff $u_{b}=\frac{1}{\epsilon}$.
@@ 2.2 Subregion volume complexity
In this section we will study the subregion complexity of the dipole deformed theory, namely the infinitely long strip: s $\frac{-W}{2}<y,z<\frac{W}{2}$ and $\frac{-l}{2}<x<\frac{l}{2}$ with $W\rightarrow\infty$. The subregion complexity will be given by the maximal volume of co-dimension one spacelike surface bounded by the co-dimension two Ryu-Takanayagi(RT) surface, homologous to the subregion Alishahiha:2015rta . The metric on co-dimension two surface,
@@ 2.3 Action Complexity
where, $\lambda$ is t’Hooft coupling. The leading term, displays a volume law scaling and is independent of the dipole-deformation parameter $a$. Next, notice that the full action complexity, coincides with the pure AdS results which matches the volume complexity upto a numerical factor, when the the t’Hooft coupling is one. This is a remarkable feature of DDSYM theory that the complexity has an explicit dependence on the ’t-Hooft coupling parameter. Perhaps this points to the validity of holographic theory even for finite t’Hooft coupling, i.e. beyond the naive supergravity regime, in contrast to the local SYM-AdS duality case. Asymptotically (large $u_{b}$),
@@
    The strong nonlocality regime result (64) is worth reflecting on for a moment. For complexity to be positive definite, the t’Hooft coupling $\lambda$ should satisfy the (upper) bound $\lambda^{2}<1+\ln{3}$. We speculate that this upper bound signals the breakdown of the UV-completeness of the DDSYM theory analogous to the LST case Katoch:2022hdf .
@@ 3 Noncommutative SYM theory: Anisotropic case
The impact of noncommutativity on the dual field theory is profound. Two-point correlation functions exhibit a characteristic UV/IR mixing — high energy (UV) behavior becomes entangled with long-distance (IR) physics, a signature of the underlying non-local interactions. Wilson loops, which in SYM serve as order parameters for confinement and phase transitions, acquire a dipole-like structure reflecting the noncommutative smearing of charge distributions. Furthermore, recent extensions of these analyses to quantum information theoretic measures, such as holographic entanglement entropy and computational complexity Karczmarek:2013xxa , reveal modified scaling laws and phase structures. Also see Fischler:2013gsa ; Couch:2017yil ; Fischler:2018kwt ; Eccles:2021zum for related work on aspects of entanglement entropy, complexity, information and entanglement spreading, and chaos and operator growth in the anisotropic NCSYM in the holographic context.
@@ 3.1 Volume complexity for the anisotropic case
$V(t_{*})$ $=\int d^{9}\sigma\penalty 10000\ e^{-2(\phi-\phi_{0})}\sqrt{G}=R^{9}\,W^{3}\,\omega_{5}\,\int_{u_{*}}^{u_{b}}du\,u^{2}\sqrt{1-u^{4}t^{\prime}(u)^{2}},$ (68)
Here $W$ is an IR cutoff in the boundary $W/2\leq x,y,z\leq W/2$. By inspection of (68), it is evident that the global maxima occurs when $t^{\prime}(u)=0$. implying that the volume complexity is
$\mathcal{C}_{V}(t_{*})$ $=\frac{\omega_{5}}{3G_{N}^{(10)}}\,W^{3}u_{b}^{3}$ (69)
Again, similar to the dipole case, the volume complexity is independent of noncommutative parameter $a$. This could have been predicted from the volume functional (68), where the noncommutative deformation factors from the metric coefficients of the two noncommutative directions, and the deformation factors arising from the dilaton profile cancels each other, even before maximization, a purely dimensional accident or conspiracy. In terms of the field theory picture, one can interpret this in the following way: the loss of degrees of freedom due to noncommutativty in the UV is exactly compensated by the usual addition of degrees of freedom as one RG-flows from IR to the UV. As we will shortly see this dimensional conspiracy is evaded when we consider the fully isotropic NCSYM theory.
@@ 3.2 Subregion Volume complexity (CV) for the anisotropic case
To compute subregion complexity we focus our attention on the portion of the maximal volume slice which is contained within the Ryu-Takayanagi (RT) surface, and which is also homologous to the boundary subregion of interest. We take the bondary subregion to be the infinitely long strip, $\frac{l}{2}<x<\frac{l}{2},\ -\frac{W}{2}<y,z<\frac{W}{2}$ in the limit $W\to\infty$ with $l$ fixed. In this case the pullback metric on the bounding codimension 2 (RT) surface is
@@
Since the above integration is not analytically tractable in general, and to gain traction we look at two different regimes, namely near AdS (large subregions) and far from AdS (small subregions) to obtain $l$ in terms of $u_{*}$, which will be useful to obtain complexity in terms of subregion length $l$ in later subsections.
@@
Similar to near AdS case, we can also work out $l$ in terms of $u_{*}$ in the far from AdS regime as,
@@
The above integral is not analytically tractable, so we plot the numerical integration results as shown in figure 2 for small values of the noncommutative deformation parameter (for large values of the deformation the numerics are not quite reliable or stable). The value of the critical subregion size (length) for three different values of $a$ extracted from the plots (in units of AdS radius) are as displayed in the table below. We also provide analytical estimates of the critical subregion size for comparison.
@@ 3.3 Action complexity (CA) for the anisotropic case (10 D)
The low energy limit of Type IIB string theory reduces to the following SUGRA bulk action Couch:2017yil with Neveu-Schwartz, Ramond-Ramond and Chern-Simmons sector outlined below
@@
  • Einstein-Hilbert term [^7] Here we use $\frac{\omega_{5}}{2\kappa^{2}_{10}}=\frac{\omega_{5}}{16\pi G_{N}^{(10)}}=\frac{1}{16\pi G_{N}^{(5)}}$
@@
The arguments of logarithm have been rendered dimensionless by rewriting them in the units of AdS radius. Moreover, $l_{ct}$ has been chosen to be one in the units of AdS radius without tempering the known universal result for the local field theory dual to AdS bulk. It can be seen that for the anisotropic case of noncommutative field theory, the action complexity is positive only in the range $\ln a<\frac{8}{3}$. Conversely positivity of action complexity imposes an upper bound on the noncommutativity parameter $a<\exp{\left(8/3\right)}$. This feature of action complexity is parallels with the action complexity of the DDSYM theory. However, unlike the DDSYM theory, but instead just like the local SYM field theory, the UV divergence scales with volume $L^{3}$! We defer our speculations about this counterintuitive result till Sec. 5.
@@ 4 Noncommutative SYM theory: Isotropic case
In the previous two cases of DDSYM and anisotropic NCSYM, the deformation had a preferred direction(s) and the full volume complexity turned out to be independent of the deformation parameter, i.e. turned out to be identical to undeformed SYM (pure AdS). We suspect this is a dimensional accident, i.e. artifact of the fact that there are only two spatial directions along which the deformation is turned on, and if the the deformation democratically involves all spatial and temporal directions, the full volume complexity will exhibit features distinct from the undeformed theory. To confirm our conjecture, in this section, we consider the Maldacena-Russo solution of a D3-brane in a constant NS $B$ field background given as Lin:2000ps ,
@@ 4.2 Action complexity (CA) for the isotropic case (5D)
Here we, as we did in the previous sections, in the end we have rescaled every argument of the logarithm in the units of AdS radius. Moreover, we have also set $l_{ct}$ to be one in the AdS radius scale without tempering the universal piece coming from the known pure AdS result. It is evident that the overall action complexity for the noncommutative case is negative unless the noncommutativity parameter satisfies the bound $\ln\left(\frac{a^{64}r_{b}^{58}}{5^{6}}\right)<4$ or $\left(\frac{a^{64}r_{b}^{58}}{5^{6}}\right)<e^{4}$. Moreover, the leading order scaling with the noncommutative coupling is the product of quadratic and logarithmic scaling with the nonlocal coupling parameter. It is however clear that due to the involvement of the UV scale the complexity cannot be made positive in any range of values in the coupling space. This is hinting at the theory being UV incomplete, i.e. it has UV instability (in past work Katoch:2023dfh we have also noted that negative holographic complexity could imply a system without a stable ground state). Also, the action complexity in the noncommutative case scales with the fifth power of the UV cutoff, unlike the local field theory that is expected to scale cubically in the UV cutoff scale. This power law departure from the extensive scaling is interpreted as the effect of nonlocality on the action complexity.
@@ 5 Discussions and Outlook
Finally, we discuss few avenues for future work. The most natural and immediate extension would be to include the effects of finite temperature and finite charge/chemical potential in complexity for nonlocal field theories using the holographic charged black hole/brane SUGRA backgrounds instead Eccles:2021zum ; Couch:2017yil . This is always an interesting exercise to carry out because the dual theory is nonlocal and it has the potential to shed light on novel physical effects associated purely with finite temperatures which might not appear in the local theories or for which we might not have any intuition from local theories. A second avenue to pursue is the subregion complexity for the isotropic NCSYM case for which the numerical estimates obtained using Mathematica or Python were unstable and unreliable for even small or moderate values of the deformation parameter $a$ - the numerical errors rapidly escalate to large values in course of the numerical solution of the RT curve. The numerics do not reproduce the pure AdS quadratic scaling of holographic entanglement entropy itself, let alone subregion complexity. It would be desirable to come up with a numerical scheme where the errors could be brought under control for moderate to large values of the deformation parameter $a$ and tie up this loose end in our present work. Another instructive, although perhaps less interesting avenue to explore will be to compute the subregion action complexity for all the cases considered in this paper. Finally, casting a wider net, one can investigate the holographic complexity (in addition to entanglement entropy) properties of more general nonlocal field theories dual to string backgrounds which are not asymptotically AdS, e.g. the nonlocal (C)FTs dual to null boundary holography Ferko:2025elh ; Dei:2025ryd ; Chakraborty:2025rfm . These are prototypes of celestial holography, and will help us explore holography/quantum gravity in backgrounds beyond the asymptotically AdS paradigm.
@@ Acknowledgements
We thank the organizers and participants of the BIRS-CMI workshop on “Quantum Gravity and Information Theory: Modern Developments” (9-14 Nov. 2025) which offered us a chance to present our results. We also thank Juan F. Pedraza for drawing our attention to the relevant works Fischler:2013gsa ; Couch:2017yil ; Eccles:2021zum . The work of SP and SR is supported by the CRG grant of the Anusandhan National Research Foundation (ANRF), Department of Science and Technology (DST), CRG/2023/001120 (“Many facets of complexity: From chaos to thermalization”). The work of SR is also supported by the IIT Hyderabad seed grant SG/IITH/F171/2016-17/SG-47 as well the IITH funds RDF/IITH/F171/SR. The work of GK is supported by IIT Indore Sponsored Research Project no. CRG/2023/000904 (“From Modular Hamiltonian to Operator Algebra in Gauge/Gravity Dictionary”).
@@ Appendix A Entanglement entropy results for dipole deformed SYM
As the integral involved in area (17) is analytically not tractable, and we proceed numerically and study the behavior of area with subregion length (18) in figure 3. This shows a phase transition in entanglement entropy. This transition has been already observed in Karczmarek:2013xxa for a near boundary (large $u$) RT surface. Here we have done a full numerical analysis and observed that as subregion length increase the entanglement entropy (area) becomes independent of subregion length as well as the deformation parameter. In the near AdS region when $au<<1$, and the leading term in entanglement entropy is same as that of pure AdS. For small sub-region length degree of freedom inside the strip will be entangled to the degree of freedom outside the strip and the entanglement entropy will be proportional to the subregion length $l$. A limiting analysis of the area agrees with this interpretation.
@@
For near AdS region (weak nonlocality) $(au<<1)$:
@@
Far from AdS region (strong nonlocality) $(au>>1)$: The entanglement entropy is given by,